Grade 11Algebra

I x-1l^[(Lnlxl)^2-Lnx^2]=lx-1l^3;find the value of x.Sir could you clear why not to equate Lnlxl^2-Lnx^2 =3?

Sigma Samantasinghar
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
Case 1: when 0

U can equate (Lnlxl)^2-Lnx^2 =3
or let logx = t
t^2 – 2t – 3 = 0
it has roots 
t = -1 and t = 3
thus,
logx = -1 and logx = 3
and
x = 1/e  
for x = e^3 solution rejected as e^3 is > 1.
 
Case 2: when x> 1
Again, t will have two roots as:
t = -1 and t = 3
x = 1/e  and x = e^3
For x > 1 1/e is rejected and x = e^3 is accepted.
 
Now intersection for both cases dnot satisfies them simulatneously.
Thus  (Lnlxl^2)-Lnx^2 =3
cannot be equated.
 
Thus it helps!